The Butler-Perkins Company (BPC) must decide between two mutually exclusive projects. Each costs $6,500 and has an expected life of 3 years. Annual project cash flows begin 1 year after the initial investment and are subject to the following probability distributions:

Project A Project B
Probability Cash Flows Probability Cash Flows
0.2 $6,000 0.2 $0  
0.6 $6,500 0.6 $6,500  
0.2 $7,000 0.2 $18,000  

BPC has decided to evaluate the riskier project at 13% and the less-risky project at 8%. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the questions below.

Open spreadsheet

  1. What is each project's expected annual cash flow? Round your answers to two decimal places.

    Project A: $  

    Project B: $  

  2. Project B's standard deviation (σB) is $5,822.37 and its coefficient of variation (CVB) is 0.78. What are the values of (σA) and (CVA)? Round your answers to two decimal places.

  3. σA = $  

    CVA =

  4. Based on the risk-adjusted NPVs, which project should BPC choose?

    _________Project AProject B

  5. If you knew that Project B's cash flows were negatively correlated with the firm's other cash flow, but Project A's cash flows were positively correlated, how might this affect the decision?

    _________This would make Project B more appealing.This would make Project B less appealing.

    If Project B's cash flows were negatively correlated with gross domestic product (GDP), while A's cash flows were positively correlated, would that influence your risk assessment?

    _________This would make Project B more appealing.This would make Project B less appealing.

B C $6,500.00 3 $500.00 A 1 Project risk analysis 2 3 Costs, Projects A and B 4 Expected life of projects (in years) 5 DifferA C D 13.00% 8.00% Formulas #N/A #N/A #N/A 19 Discount rate, risky project 20 Discount rate, less risky project 21 22 Calcula#N/A 37 Calculation of Risk-Adjusted NPVs: 38 NPVA 39 NPVB 40 Which project should be chosen? 41 #N/A #N/A
Answer #1

1.
=0.2*6000+0.6*6500+0.2*7000
=6500.00

2.
=0.2*0+0.6*6500+0.2*18000
=7500.00

3.
=sqrt(0.2*(6000-6500.00)^2+0.6*(6500-6500.00)^2+0.2*(7000-6500.00)^2)
=316.22777

4.
=316.22777/6500.00
=0.04865

5.
NPV of A=-6500+6500/1.08+6500/1.08^2+6500/1.08^3=10251.13042

NPV of B=-6500+7500/1.13+7500/1.13^2+7500/1.13^3=11208.64448

Choose Project B

6.
This would make Project B more appealing.

7.
This would make Project B more appealing.